A newspaper, USA- yesterday, prints the same number of pages every day. Readers like the news and photographs in the paper, but dislike advertising. As a result, as the number of pages devoted to ads increases, the number of readers declines. This relationship is described by the equation R=120-9A, where R is the number of readers (in thousands) and A is the number of pages of advertising. At what level of A does no one buy the newspaper?
Answer by stanbon(48568)

This relationship is described by the equation R=120-9A, where R is the number of readers (in thousands) and A is the number of pages of advertising. At what level of A does no one buy the newspaper?
----------------
Ans: When R = 0.
0 = 120-9A
9A = 120
A = 40/3
A = = 13 1/3 pages of advertising
====================================
Cheers,
Stan H.